OpenStudy (anonymous):

You own 7 books and are taking 5, how many ways can you choose which books to take? Answer is 21, how do I get there?

OpenStudy (cwrw238):

this is the number of combination of 5 from 7 which is the same as number combinations of 2 (7-5) from 7 and is 7*6 --- = 21 2* 2

OpenStudy (cwrw238):

sorry the denominator is 2*1

OpenStudy (anonymous):

K, I was going 5 4 3 2 1 Thanks.

OpenStudy (anonymous):

nPr = \[n! \over (n - r)!\] \[7! \over (7 - 5)!\] \[7! \over 2!\] \[7 * 6*5*4*3 \]

OpenStudy (cwrw238):

the maths textbooks give the formula n! ------ (n-r)! r! which simplifies to the way i stated

OpenStudy (anonymous):

2520 :0

OpenStudy (anonymous):

Thanks to everyone, really. :)

OpenStudy (cwrw238):

= 7*6*5*4*3*2*1 7*6 ----------- = --- 5*4*3*2*1 * 2* 1 2*1

OpenStudy (anonymous):

@cwrw238 Isn't this a permutation because the books aren't being replaced?

OpenStudy (cwrw238):

hhmm - not sure to be honest - now you mention it

OpenStudy (anonymous):

Yeah. I think it is. In that case the formula is \[n! \over (n - r)!\]